A siren mounted on the roof of a firehouse emits sound at a frequency of 900Hz. A steady wind is blowing with a speed of 15.0m/s. Taking the speed of sound in calm air to be 343 m/s , find the wavelength of the sound

(a) upwind of the siren

(b) downwind of the siren.

Firefighters are approaching the siren from various directions at 15.0m/s


1
Expert's answer
2021-02-24T12:51:12-0500

(a)

"\lambda=\dfrac{v_s-v_{wind}}{f}=\dfrac{343\ \dfrac{m}{s}-15\ \dfrac{m}{s}}{900\ Hz}=0.364\ m."

(b)

"\lambda=\dfrac{v_s+v_{wind}}{f}=\dfrac{343\ \dfrac{m}{s}+15\ \dfrac{m}{s}}{900\ Hz}=0.398\ m."